# Where to find Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

### v2

@@ A.1 Model Setup and Assumptions
A.2 Routing Kernel Definition
@@
A.3 Main Theorem
Theorem A.1 (Generalization Bound for MoE Experts).
@@ A.3 Main Theorem
A.4 Connecting Edit Distance to Effective Dimension
@@
Thus the bound shows that simple (low-dimensional) routing kernels generalize better. We now empirically test how routing pattern complexity (via edit distance) correlates with this effective dimension $d_{\rm eff}=\text{Tr}(\bm{H}^{*}(\bm{H}^{*}+\lambda I)^{-1})$.
@@
Our setup involves a lightweight MoE architecture with 16 experts and a 100-dimensional input space. Input data are synthetic Gaussian clusters, providing a controlled environment to analyze pre-defined routing strategies. For each experimental trial, 200 samples are split into training and validation sets. We train a one-layer MoE model with ReLU activations where only expert networks are trainable; the router’s weights are initialized once and then frozen. To explore diverse fixed routing patterns, we perform multiple independent runs, each starting with a different random initialization of the fixed router.
@@
Following training, we extract the average edit distance of the pre-determined expert routing paths and the effective dimension. As illustrated in Figure A.1, we observe a strong positive correlation between these two quantities. This empirically demonstrates that the structural properties of fixed routing, measured by edit distance, meaningfully impact the complexity of the function space learned by the experts, as predicted by $d_{\rm eff}$. This alignment supports edit distance as a valuable diagnostic for MoE generalization.
@@
Figure A.1: Correlation of Edit Distance with Effective Dimension. Each point represents an independent experimental run where the router is initialized with a different random seed and then fixed for expert training. This varying initialization leads to diverse fixed routing patterns. We observe a strong positive correlation. This suggests that the structural diversity of pre-defined routing significantly impacts the complexity of the function space learned by experts.
Appendix B Dataset Details
We conduct our evaluation on four domains used during OLMoE pretraining: math, code, commonsense, and domain-specific (scientific text). For each domain, we construct two sets:
  • Training set: 10,000 examples randomly sampled from the corresponding OLMoE pretraining data.
  • Test set: 10,000 examples from a distinct dataset in the same domain, not used during pretraining.
In the domain-specific setting, we further filter the pretraining set to include only samples containing the phrase “Organic Chemistry” to create a more focused training distribution.
Table A.1 summarizes the source links for all datasets.
Table A.1: Sources of training and test data used in our domain-level evaluation. All datasets accessed via HuggingFace.
Domain Pretraining Source Test Dataset
Math open-web-math/open-web-math hkust-nlp/dart-math-pool-math
Code bigcode/starcoderdata evalplus/humanevalplus
Commonsense emozilla/dolma-v1_7-books tau/commonsense_qa
Domain-Specific blester125/mediawiki-dolma (filtered by “Organic Chemistry”) cais/mmlu ("college_chemistry" subset)
In addition to the training and test sets used for evaluation, we compute validation loss on separate domain-specific datasets to monitor in-domain learning dynamics and assess their relationship to generalization. Specifically, we aim to quantify the correlation between validation loss and test accuracy, to test whether validation loss can serve as a predictor of generalization in each domain.
To this end, we select distinct validation datasets for each domain, ensuring they are disjoint from both the training and test sets while remaining representative:
Table A.2: Datasets used for computing validation loss, used to assess correlation with test accuracy.
Domain Validation Dataset
Math qwedsacf/competition_math
Code newfacade/LeetCodeDataset
Commonsense kowndinya23/flan2021-commonsense
Domain-Specific chemNLP/chemistry-bookshelves-merged
These datasets are used solely for loss computation and correlation analysis.
Appendix C Experiment Details
Instruction Tuning. The raw OLMoE checkpoints exhibit weak instruction-following ability, making direct evaluation of generalization behavior unreliable—particularly in text-based domains such as commonsense reasoning or code completion. To address this, we apply light instruction tuning using LoRA [6], not to improve task performance per se, but to make the pretrained model follow prompts in a human-readable way. This ensures we can evaluate generalization while preserving the original pretraining behavior.
  • Math: We finetune each checkpoint using the same math SFT dataset meta-math/MetaMathQA employed by OLMoE for 3 epochs with LoRA.
  • Code: We use the HuggingFaceH4/CodeAlpaca_20K dataset and finetune for 3 epochs.
  • Commonsense & Domain-Specific: We reuse the instruction-tuned checkpoint trained on yahma/alpaca-cleaned, shared across both domains.
Key configurations of LoRA include:
  • Target modules: q_proj, k_proj, v_proj
  • LoRA rank: 32
  • LoRA dropout: 0.1
  • Learning rate: 5e-5
  • Batch size: 4
  • Max sequence length: 2048
This lightweight finetuning ensures reliable promptability while preserving the pretrained model’s generalization behavior.
Applying Min-K%++ for Training-Test Separation. To maintain a rigorous separation between the training and test datasets, we employ the Min-K%++ membership inference attack [26], a state-of-the-art, reference-free method designed to detect whether specific data samples are part of a model’s training corpus.
In our experiments, we apply the Min-K%++ method to the test dataset to identify samples that the model is most likely to have encountered during training. Determining an optimal threshold for classification is challenging due to variability in model behaviors and data distributions. To address this, we adopt a conservative approach by removing the top 10% of test samples with the highest Min-K%++ scores, thereby minimizing the risk of training data contamination in the test set.

### v3

@@ C.1 Model Setup and Assumptions
C.2 Routing Kernel Definition
@@
C.3 Main Theorem
Theorem C.1 (Generalization Bound for MoE Experts).
@@ C.3 Main Theorem
C.4 Connecting Edit Distance to Effective Dimension
@@
The bound shows that simple (low‐dimensional) routing kernels generalize better. We now empirically test how routing pattern complexity (via edit distance) correlates with this effective dimension $d_{\rm eff}=\text{Tr}(\bm{H}^{*}(\bm{H}^{*}+\lambda I)^{-1})$.
@@
Our setup involves a lightweight MoE architecture with 16 experts and a 100-dimensional input space. Input data are synthetic Gaussian clusters, providing a controlled environment to analyze pre-defined routing strategies. For each experimental trial, 200 samples are split into training and validation sets. We train a one-layer MoE model with ReLU activations where only expert networks are trainable; the router’s weights are initialized once and then frozen. To explore diverse fixed routing patterns, we perform multiple independent runs, each starting with a different random initialization of the router.
@@
Following training, we extract the average edit distance of the pre-determined expert routing paths and the effective dimension. As illustrated in Figure 8, we observe a strong positive correlation between these two quantities. This empirically demonstrates that the structural properties of fixed routing, measured by edit distance, meaningfully impact the complexity of the function space learned by the experts, as predicted by $d_{\rm eff}$. This alignment supports edit distance as a valuable diagnostic for MoE generalization.
@@
Figure 8: Correlation of Edit Distance with Effective Dimension. Each point represents an independent experimental run where the router is initialized with a different random seed and then fixed for expert training. This varying initialization leads to diverse fixed routing patterns. We observe a strong positive correlation. This suggests that the structural diversity of pre-defined routing significantly impacts the complexity of the function space learned by experts.
Appendix D Routing Dynamics Analysis
To investigate whether routing evolution arises primarily from router-layer updates or other weight updates, we analyze convergence speed and stability across parameter types.
Results.
Table 3 reports convergence speed (magnitude drop-off over time) and stability (consistency of updates). Router weights adapt fastest while expert weights stabilize most gradually, indicating a division of labor: routing decisions are steered by fast-updating routers atop slow-evolving expert features. Attention layers lie in between.
Table 3: Convergence speed and stability across parameter types. Higher values indicate faster convergence or greater stability.
Parameter Type Convergence Speed $\uparrow$ Stability $\uparrow$
Router 0.151 (fastest) 5.63
Expert 0.063 (slowest) 7.32
Attention 0.083 6.01
Layerwise Routing Dynamics.
To localize routing evolution, we compute the edit distance of routing assignments across layers. Layers 0–3 exhibit the most post-convergence change, suggesting that early layers remain active in refining pathways.
Metric Computation.
We periodically snapshot parameter L2 norms during training. For each parameter type, we compute the mean absolute norm change between checkpoints, normalized by average scale. Convergence speed is defined as the drop in change magnitude from early to late training, while stability is the inverse standard deviation of these normalized changes. This simple but robust analysis consistently reveals router–expert asymmetries.